OpenStudy (sphott51):

solve log(x+9) -logx = 3 I know the - is division so, x/ (x+9)

OpenStudy (anonymous):

base ten right?

OpenStudy (sphott51):

Yeah I thinks so

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \log(x+9) -\log x = 3 }\) \(\large\color{black}{ \displaystyle \log\left(\frac{x+9}{ x}\right) = 3 }\) Yes, that is right

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \log_{10}\left(\frac{x+9}{ x}\right) = 3\quad \Rightarrow \quad \left(\frac{x+9}{ x}\right)=10^3}\)

OpenStudy (sphott51):

And 10^3 is 1000

OpenStudy (solomonzelman):

yes

OpenStudy (sphott51):

So what do I do next?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{x+9}{ x}=1000 }\) solve for x.

OpenStudy (solomonzelman):

Just using algebra, you can solve for x easily.

OpenStudy (sphott51):

1000/9 = 111.111 so its just 111?

OpenStudy (anonymous):

no

OpenStudy (sphott51):

?